enlo Company distributes a single product. The company’s sales and expenses for last month follow:

Total    Per Unit
  Sales $ 624,000 $ 40     
  Variable expenses 436,800 28     
  Contribution margin 187,200 $ 12     
  Fixed expenses 154,800
  Net operating income $   32,400

Required:

1. What is the monthly break-even point in unit sales and in dollar sales?

2. Without resorting to computations, what is the total contribution margin at the break-even point?

3-a. How many units would have to be sold each month to earn a target profit of $61,200? Use the formula method.

3-b. Verify your answer by preparing a contribution format income statement at the target sales level.

4. Refer to part 3 and now assume that the tax rate is 30%. How many units would need to be sold each month to an after-tax target profit of $61,200?  (Round the final answer to the nearest whole number.)

5. Refer to the original data. Compute the company's margin of safety in both dollar and percentage terms. (Round your percentage answer to 2 decimal places (i.e .1234 should be entered as 12.34).)

6. What is the company’s CM ratio? If monthly sales increase by $71,000 and there is no change in fixed expenses, by how much would you expect monthly net operating income to increase?
Answer #1

1) Break even unit = 154800/12 = 12900 Units

Break even sales = 12900*40 = $516000

2) total contribution margin = $154800

3a) Required unit = (154800+61200)/12 = 18000 Units

3b) Contribution margin income statement

Sales (18000*40) 720000
Variable cost (18000*28) 504000
Contribution margin 216000
Fixed cost 154800
Net income 61200

4) Income before tax = 61200*100/70 = 87429

Required unit = (154800+87429)/12 = 20186 Units

5) Margin of safety = 624000-516000 = 108000

Margin of safety percentage = 108000/624000 = 17.31%

6) CM ratio = 12/40 = 30%

Net operating income increase = 71000*30% = 21300